The Cost of Capital: Cost of New Common Stock If a firm plans to issue new stock, flotation costs (investment bankers' fees) should not be ignored. There are two approaches to use to account for flotation costs. The first approach is to add the sum of flotation costs for the debt, preferred, and common stock and add them to the initial investment cost. Because the investment cost is increased, the project's expected return is reduced so it may not meet the firm's hurdle rate for acceptance of the project. The second approach involves adjusting the cost of common equity as follows: The difference between the flotation-adjusted cost of equity and the cost of equity calculated without the flotation adjustment represents the flotation cost adjustment. Quantitative Problem: Barton Industries expects next year's annual dividend, D1, to be $1.60 and it expects dividends to grow at a constant rate g = 4.5%. The firm's current common stock price, P0, is $20.50. If it needs to issue new common stock, the firm will encounter a 4.8% flotation cost, F. Assume that the cost of equity calculated without the flotation adjustment is 12% and the cost of old common equity is 11.5%.

What is the flotation cost adjustment that must be added to its cost of retained earnings? Round your answer to 2 decimal places. Do not round intermediate calculations. ________%

What is the cost of new common equity considering the estimate made from the three estimation methodologies? Round your answer to 2 decimal places. Do not round intermediate calculations.

________%
Answer #1

cost of equity with flotation cost adjustment = [Dividend next year/current share price*(1-flotation cost)] + constant dividend growth rate

cost of equity with flotation cost adjustment = [$1.60/$20.50*(1-0.048)] + 0.045 = [$1.60/($20.50*0.952)] + 0.045 = ($1.60/$19.516) + 0.045 = 0.0819840131174421 + 0.045 = 0.1270 or 12.70%

flotation cost adjustment = cost of equity with flotation cost adjustment - cost of equity without the flotation adjustment = 12.70% - 12% = 0.70%

cost of new common equity = cost of old common equity + flotation cost adjustment = 11.5% + 0.70% = 12.20%